OpenStudy (anonymous):

Working together, Tim and Jeff can detail a car in 2.5 hours. Working alone, Tim can detail the same car in 4.5 hours. Approximately how long will it take Jeff to detail the car if he is working alone?

OpenStudy (anonymous):

Time together Time together ------------ + ----------- = 1 Time alone Time alone Generic equation

OpenStudy (anonymous):

2.5 2.5 -- + --- = 1 4.5 x Multiply by common denominator of 4.5x

OpenStudy (anonymous):

4.5x(2.5) 4.5x(2.5) -------- + -------- = 4.5x(1) 4.5 x

OpenStudy (anonymous):

Cancel the denominators into the numerators

OpenStudy (anonymous):

x(2.5) + 4.5(2.5) = 4.5x(1)

OpenStudy (anonymous):

2.5x + 11.25 = 4.5x

OpenStudy (anonymous):

Solve. 2.5x + 11.25 = 4.5x 11.25 = 2x 5.625 hours

OpenStudy (anonymous):

wow..thanks.